# ITCMA: A Generative Agent Based on a Computational Consciousness Structure

### v1

@@ 3.1.2. Stream of Consciousness and Internal Time-Consciousness
3.1.3. Memory and Protention
Accidental memory reproduction involves involuntary memory and is built on the similarity between two different sensations at different times. Therefore, the current primal impression is not solely determined by its previous retention. In the current primal impression, a certain memory from the past can be awakened. If it has been awakened at this moment, it is fused into the current consciousness. Thus, equation (3) can be specifically expressed as follows:
@@ 3.1.3. Memory and Protention
Before providing the $M$ algorithm, we must first provide the function $FieldSim(f^{x},f^{y})$ to calculate the similarity between two different $f$ values. Because $f$ is an $m\texttimes(n+3)$ matrix with an indeterminate number of rows $m$, to compute the degree of difference between the two, for a $f^{x}$ with a shape of $a\texttimes(n+3)$ and an $f^{y}$ with a shape of $b\texttimes(n+3)$, the degree of difference between the two is computed using $FieldSim(f^{x},f^{y})$ as follows:
(7) $N_{m}=\begin{bmatrix}{N^{1}}_{m}&\dots&{N^{n}}_{m}\end{bmatrix}$
(8) $pos_{m}=\begin{bmatrix}\theta_{m}&\varphi_{m}&\gamma_{m}\end{bmatrix}$
@@
The function $lev_{Re^{t}+PI^{t},Memory^{i,j}}(i,j)$ provides the Levenshtein distance between $\begin{bmatrix}Re^{t}&PI^{t}\end{bmatrix}$ and $Memory^{i,j}$. Due to the structural differences between the field string and the alphabetic string, the calculation method for their Levenshtein distance needs to be modified. Specifically, we did not set the number of steps required for each element change to be 1 but to the difference $diff(f^{x},f^{y})$ between the two elements modified by this change as follows:
@@
(15) $A=\begin{bmatrix}{f^{a}}_{1}&{f^{a}}_{2}&\dots&{f^{a}}_{n}\end{bmatrix}$
(16) $B=\begin{bmatrix}{f^{b}}_{1}&{f^{b}}_{2}&\dots&{f^{b}}_{m}\end{bmatrix}$
@@
When memory comes to this moment, it becomes working memory and is used for guidance regarding the present; it is used, along with retention and primal impression, to predict protention. However, an agent is not just a prediction algorithm. Heidegger [42] regarded plans toward the future as a more fundamental time dimension, thus forming a temporal structure with the future as the priority dimension. This implies that the prediction of protention is not without a trend or is only based on previous trends. There is a driving force $d^{t}$ in its prediction at time $t$.
3.1.4. Drive and Emotions
According to Held [43], the source of this drive is “hope.” The hope expressed here is subsumed under the superordinate concept of “expectation.” We believe that the role of expectation is associated with future feelings, and this association is infected by specific emotions. We first expect something and then assign a certain degree of emotion to this expectation. Therefore, $d^{t}$ can be expressed as follows:
@@ 3.1.4. Drive and Emotions
that is, the previous moment’s drive $d^{t-1}$ plus a bias consisting of the Hadamard product of the current emotions and their weights. In the equation, $E$ is the vector of emotions, while $W$ is the dynamic weight matrix. The consideration of emotions as a driving force here is due to the forced updating of the perceptual field and the prediction of the phenomenal field, which leaves room for the ITCM to introduce emotions. For example, the special feature of the emotion fear is that it is unknown and uncontrollable, which is highly similar to the prediction and perception of objects. The “unknown and uncontrollable” mentioned here corresponds precisely to the dimension of dominance. In addition, for the arousal and pleasure of emotions, we can temporarily assume that the former comes from the amplitude of changes in the monad flow, which is related to passive attention [44], and that the latter comes from satisfying survival needs, such as eating [45]—of course, the actual situation is much more complex.
In the PAD model, emotions can be described and distinguished through three dimensions: pleasure, arousal, and dominance [47; 48; 49]. Of course, hope as a feeling is an indicator of emotion. It is not another short-term impulse but constantly impacts from its own dimension into the action level. Therefore, such a feeling can also be interpreted as a continuous dimensional model:
(19) $E=\begin{bmatrix}P&A&D\end{bmatrix}$
(20) $W=\begin{bmatrix}\omega_{P}&\omega_{A}&\omega_{D}\end{bmatrix},$
where $A$ represents the arousal dimension. Based on the mechanism of passive attention, it is modeled here as the degree of change from elements in retention to primal impression. $D$ represents the dominance dimension. Based on its original meaning (degree of control over the environment), it represents the difference between the predicted protention $Pro^{t-1}$ at the previous moment and the primal impression $PI^{t}$ at this moment. $P$ represents the pleasure dimension, which quantifies the degree of satisfaction of the agent’s desire and the degree of avoidance of pain. The calculation of $desire\in[0,\infty)$ and $pain\in[0,\infty)$ is defined according to specific situations. For example, in reinforcement learning tasks, something desired can be defined as a reward, while pain can be defined as a punishment. Specifically, there are the following:
@@
3.2. ITCM-Based Agent
Park et al. [3] proposed the original generative agent structure in which all content is recorded and inferred in natural language, enabling the structure to utilize an LLM. Influenced by this, Xi et al. [4] proposed a general conceptual framework for generative agents that includes three key components: the brain, perception, and action. The structure of the ITCM-based agent (ITCMA) is similar in that perception serves as input from the environment, and action serves as output to the environment. After receiving action, the environment forces the agent to form a perception as a re-action to it. Reaction is the carrier of this action—the brain is not equivalent to the LLM here but is replaced by the complete ITCMA, and the LLM is only used as a tool. The structure is shown in Figure 3.
Figure 3. Conceptual framework of the ITCMA.
Figure 3 is based on the specific ITCM structure, which can be further refined. The perception provided can be considered the primal impression $PI^{t}$ at $t$. After receiving $C^{t-1}$ of time $t-1$ to form a consciousness channel $C^{t}$, the agent extracts the activated memory $AM$ and its subsequent content from long-term memory, thus forming the protention content $Pro^{t}$ through TSFM. It is integrated with the driving force $d^{t}$ into a natural language format and provided to the LLM to form an action, as shown in Figure 4.
Figure 4. Detailed structure of the ITCMA.
Maslow argues that there are two types of human behavior: coping behaviors and expressive behaviors [80]. We can understand expressive behaviors as mechanical reflexes, such as finger lifting, and coping behaviors as behaviors in a broader sense, which include running and eating. Bobick [79] stated that human movement can be divided into movement, activity, and action, while He [81] suggested that human movement can be divided into behavior, action, and basic action. They both described action as the basic unit of human movement and that action constitutes more complex actions and behaviors. For example, a finger-bending action is part of grabbing; grabbing, along with other actions, constitutes eating behavior.
Therefore, the actions obtained in ITCMA should be considered as coping behaviors, and when deployed to an agent that requires specific movements to manipulate, the actions that serve as coping behaviors can be decomposed into a set of learned expressive behaviors. Such movements can come from motion algorithms in robotics and can be considered as ”unconscious actions,” just like humans do not pay attention to how their limbs move when performing ”go to somewhere” actions.
Because the ITCM constructs an agent that simulates consciousness without social experience, the use of a visual language model (VLM) to aid in modeling is an option here. In complex task environments, the information captured by vision is not only the spatial topology of an object but also includes a holistic understanding of it. Using the VLM here to characterize the visual image and add it to the ITCMA workflow can somewhat bypass its early learning and growth process. However, due to the enhancement of the VLM’s unique visual ability, its language generation ability is slightly insufficient compared to an LLM of the same scale. Therefore, it is only used as a descriptive tool here and cannot replace the original LLM used to generate actions. Appendix B includes the specific implementation details and examples of the ITCMA.
@@ 4.3. Evaluation Procedure
Due to the need for fine-tuning and considerations for testing the generalizability of the model, we chose ChatGLM3-6B [56; 24] as the TSFM for generating protention. By fine-tuning and using function tools, it can generate text for the protention based on the provided information and convert it into a field format for storage. In addition, we chose GPT-4 as the LLM for generating actions.
@@ 4.4. Results
Table 1. The completion rates of different models in each evaluation segment (see set and unseen set) in Alfworld (%).
@@ 5.1. Discussion on Experimental Results
Using the proposed ITCM computational consciousness structure, this study enhances the behavioral performance of agents by simulating a part of human consciousness. This structure enhances the behavioral ability of the LLM, without which the LLM would be limited to generating text-based questions and answers and general information. Obviously, in traditional research on generative agents, the approach of equating the LLM to the brain [4] is questionable. By treating LLM only as the final decision-making part of a stream of consciousness, it can present almost human-like behavioral logic. Therefore, throughout the entire research process, we can see that the ITCMA can make inferences and decisions based on environmental conditions. As shown in Section 4.4, even an untrained ITCMA can quickly start tasks after exploring the environment.
Anderson believes that intelligent organisms have three ways to solve problems: backup avoidance, difference reduction, and means–ends analysis [58]. Among them, means–ends analysis is considered to have shown significant progress compared to the previous two and is believed to be a significant reason for the intelligence gap between humans and other organisms [59]. CoT [6] originally attempted to solve the problem of means–ends analysis, but we can see that even GPT-4 using CoT has difficulty achieving a high rate of task completion. On this basis, it can be considered that the ITCM structure provides an LLM with a better and more human-like thinking logic. This may be the basis for the generation of artificial general intelligence (AGI).
@@ 5.2. Model That Is Not Entirely Devoted to One Task
The researcher is not required to build a model from scratch each time and does not have to use one as a migration substrate for a class of tasks, such as a pre-training model (the role of a pre-training model between different classes of tasks is very limited). The researcher only has to design a mediator for different tasks that can guide the agent in generating the spatio-temporal similarity between other untrained tasks and will be written into the memory network. A model that is not entirely devoted to one task requires only one agent, and its experience existing in the memory network on the original task (even if not of the same type as the current task) can be a basis for the current task and a basis for transferability when the memory network has a certain complexity. Such a model is expected to become the foundation of Artificial General Intelligence (AGI). We have selected several relatively noteworthy points in consciousness research for analysis and discussion, as shown in Appendix C.
@@ 6. Conclusion
In this work, we proposed a generative agent framework based on a computational consciousness structure that uses consciousness channels, protention, memory, and drive to assist in implementing a specific agent. Our agent framework is designed to combine LLMs with mechanisms for synthesizing and retrieving relevant information, thereby providing formatted language model outputs. In our experiment, we deployed an ITCMA in text-based life scenarios and real-world scenarios, both of which effectively handled various tasks and sub-objectives. Our agent performed better than the baseline in the experimental scenarios, demonstrating better utility and the ability to continually evolve to target tasks. In addition, because our framework can be adapted to a specific task without training, it exhibits better flexibility and generalization capabilities. Our results indicate that the ITCMA framework has advantages and further potential in handling complex tasks and real-world scenarios.
@@ Appendix A Phenomenology for the Analysis of the Consciousness Structure
Appendix B Specific implementation details and examples of ITCMA

### v2

@@ 3.2.2. Consciousness Channel Module
The components of retention, primal impression, and protention are all $f$-representations. This structure is shown in Figure 4.
Figure 4. Consciousness channel composed of retention, primal impression, and protention. The primal impression at each moment continuously enters into the retention of the next moment, while new primal impressions and protentions are continuously perceived and generated.
3.2.3. Memory Module
A continuous consciousness stream channel is not sufficient for making accurate predictions about the future. As mentioned earlier, the memory that differs from retention needs to be taken into account. Proust [42] proposed involuntary memory, arguing that such memories cannot emerge by will. This association stems from topological similarity. When completing a mechanical task, predictions for the future should come from the experience of changes to objects in similar environments, and this experience should also be considered in migratory work. For example, when transferring the experience of playing table tennis to playing baseball, if the features are extracted, they can both be seen as ”contacting a high-speed moving object with a controllable object in the phenomenal field.” This connection is established in the high-dimensional space of memory and is also computable.
The involuntary memory is built on the similarity between two different sensations at different times. Therefore, the current primal impression is not solely determined by its previous retention. In the current primal impression, a certain memory from the past can be awakened. If it has been awakened at this moment, it is fused into the current consciousness. Thus, formula (2) can be specifically expressed as follows:
@@ 3.2.3. Memory Module
Before providing the $M$ algorithm, we must first provide the function $FieldSim(f^{x},f^{y})$ to calculate the similarity between two different $f$ values. Because a $f$ containing m objects is an $m\texttimes(n+3)$ matrix, to compute the degree of difference between the two, for a $f^{x}$ with a shape of $a\texttimes(n+3)$ and an $f^{y}$ with a shape of $b\texttimes(n+3)$, the degree of difference between the two is computed using $FieldSim(f^{x},f^{y})$ as follows:
@@
The function $lev_{Re^{t}+PI^{t},Memory^{i,j}}(i,j)$ provides the Levenshtein distance between $\begin{bmatrix}Re^{t}&PI^{t}\end{bmatrix}$ and $Memory^{i,j}$. Due to the structural differences between the field string and the alphabetic string, the calculation method for their Levenshtein distance needs to be modified. Specifically, we did not set the number of steps required for each element change to be 1 but to the difference $diff(f^{x},f^{y})$ between the two elements. Therefore, for two field strings $A\leftarrow\begin{bmatrix}{f^{a}}_{1}&{f^{a}}_{2}&\dots&{f^{a}}_{n}\end{bmatrix}$ and $B\leftarrow\begin{bmatrix}{f^{b}}_{1}&{f^{b}}_{2}&\dots&{f^{b}}_{m}\end{bmatrix}$, their Levenshtein distance $lev_{A,B}(i,j)$ is as follows:
@@
When memory comes to this moment, it becomes working memory and is used for guidance regarding the present; it is used, along with retention and primal impression, to predict protention.
3.2.4. Drive Module
The prediction of protention is not without a trend or is only based on previous trends. There is a driving force in its prediction at time $t$. We believe that such a trend is associated with future feelings, and this association is infected by specific emotions. We first expect something and then assign a certain degree of emotion to this expectation. Therefore, driving force $d^{t}$ can be expressed as follows:
@@ 3.2.4. Drive Module
that is, the previous moment’s drive $d^{t-1}$ plus a bias consisting of the Hadamard product of the current emotions and their weights. We use the PAD model as a quantification of emotions (Mehrabian, 1995, 1996; Mehrabian et al., 1997). In the PAD model, emotions can be described and distinguished through three dimensions: pleasure, arousal, and dominance. In the formula (13), $E\leftarrow\begin{bmatrix}P&A&D\end{bmatrix}$ is the vector of emotions, while $W=\begin{bmatrix}\omega_{P}&\omega_{A}&\omega_{D}\end{bmatrix}$ is the dynamic weight matrix.
The $A$ represents the arousal dimension. Based on the mechanism of passive attention, it is modeled here as the degree of change from elements in retention to primal impression. The $D$ represents the dominance dimension. Based on its original meaning (degree of control over the environment), it represents the difference between the predicted protention $Pro^{t-1}$ at the previous moment and the primal impression $PI^{t}$ at this moment. The $P$ represents the pleasure dimension, which quantifies the degree of satisfaction of the agent’s desire and the degree of avoidance of pain. The calculation of $desire\in[0,\infty)$ and $pain\in[0,\infty)$ is defined according to specific situations. For example, in reinforcement learning tasks, something desired can be defined as a reward, while pain can be defined as a punishment. Specifically, there are the following:
@@ 4.1.2. The Quadruped Robot in the Real World
It should be noted that we are not considering the joint movement algorithms of the robot here, but only controlling the robot to execute command-like actions through the Tinymal interface. Bobick [81] stated that human movement can be divided into movement, activity, and action. Another study described action as the basic unit of human movement and that action constitutes more complex actions and behaviors [83]. For example, a finger-bending action is part of grabbing; grabbing, along with other actions, constitutes eating behavior. Therefore, The actions provided by ITCMA should not be considered as movements, and when deployed to an agent that requires specific movements to manipulate, the actions can be decomposed into a set of learned motion algorithm. Such movements can be considered as ”unconscious actions,” just like humans do not pay attention to how their limbs move when performing ”go to somewhere” actions.
@@ 4.3. Evaluation Procedure
Due to the need for fine-tuning and considerations for testing the generalizability of the model, we chose ChatGLM3-6B [58; 25] as the TSFM for generating protention. By fine-tuning and using function tools, it can generate text for the protention based on the provided information and convert it into a field format for storage. In addition, we chose GPT-4 as the LLM for generating actions.
@@
To further validate the effectiveness of ITCMA and understand the interactions between its modules, we conducted an ablation study. There are three ablation architectures:
  1. 1.
    No Consciousness Channel. This architecture cannot access the contents of the consciousness channel module, including retention and protention.
  2. 2.
    No Memory. This architecture cannot access the contents of the memory module, meaning that retrieved memories are not provided to the LLM.
  3. 3.
    No Driving Force. This architecture cannot access the contents of the driving force module, meaning that emotions and further action selection tendencies are not computed, and no selection weights are assigned to the action space.
We randomly selected 10 different Alfworld environments to evaluate both the full architecture and the ablation architectures. All architectures were evaluated without training.
@@ 4.4. Results
Table 1. The completion rates of different models in each evaluation segment (seen set and unseen set) in Alfworld (%).
@@
The full ITCMA architecture achieved the best task completion rate (91$\pm$7). Under ablation conditions, the performance of the architecture declined with the removal of each module: the ablation architecture that could not access the driving force module was the next best (60$\pm$16), followed by the ablation architecture that could not access the memory module (40$\pm$15). The ablation architecture that could not access the consciousness channel module performed the worst under all conditions (30$\pm$16). The results are shown in Figure 7.
Figure 7. Comparison of different modules turned on (%).
@@ 5.1. Discussion on Experimental Results
Using the proposed ITCM computational consciousness structure, this study enhances the behavioral performance of agents by simulating a part of human consciousness. This structure enhances the behavioral ability of the LLM, without which the LLM would be limited to generating text-based questions and answers and general information. Obviously, in traditional research on generative agents, the approach of equating the LLM to the brain [4] is questionable. By treating the LLM only as the final decision-making part of a stream of consciousness, it can exhibit human-like consciousness processes through the rules provided by ITCM. Evaluation results indicate that in the Alfworld environment, the trained ITCMA can present human-like behavioral logic in the seen set. Therefore, throughout the entire research process, we can see that the ITCMA can make inferences and decisions based on environmental conditions. As shown in Section 4.4, even an untrained ITCMA can quickly start tasks after exploring the environment.
The results of the ablation study indicate that the full ITCMA architecture produces more credible actions than the ablation architectures. Each additional ablation decreases the architecture’s performance. We observed the actions of agents in the ablation study and found that the different ablation architectures exhibited distinct action patterns. The action of the ablation architecture that could not access the consciousness channel module lacked regularity and purposeful exploration, wandering aimlessly between locations, including those irrelevant to the tasks. After obtaining an item, the ablation architecture without access to the memory module will cycle between putting it down and picking it up, or moving back and forth between the location associated with the task and other locations.The ablation architecture that could not access the driving force module exhibited chaotic decision-making and made redundant plans, such as continuing to explore other locations even when already holding the required item.
Due to ITCMA’s linear computation, each module somewhat relies on the outputs of other modules. However, we believe these results still demonstrate how each module contributes to the overall effectiveness of the full architecture. For example, the consciousness channel module provides a continuous stream of action, the memory module supplies planning, and the driving force module identifies the actions most worth executing.
From other perspectives, Anderson [60] believes that intelligent organisms have three ways to solve problems: backup avoidance, difference reduction, and means–ends analysis. Among them, means–ends analysis is considered to have shown significant progress compared to the previous two and is believed to be a significant reason for the intelligence gap between humans and other organisms [61]. CoT [7] originally attempted to solve the problem of means–ends analysis, but we can see that even GPT-4 using CoT has difficulty achieving a high rate of task completion. On this basis, it can be considered that the ITCM structure provides an LLM with a better and more human-like thinking logic. This may be the basis for the generation of artificial general intelligence (AGI).
@@ 5.2. Model That Is Not Entirely Devoted to One Task
The researcher is not required to build a model from scratch each time and does not have to use one as a migration substrate for a class of tasks, such as a pre-training model (the role of a pre-training model between different classes of tasks is very limited). The researcher only has to design a mediator for different tasks that can guide the agent in generating the spatio-temporal similarity between other untrained tasks and will be written into the memory network. A model that is not entirely devoted to one task requires only one agent, and its experience existing in the memory network on the original task (even if not of the same type as the current task) can be a basis for the current task and a basis for transferability when the memory network has a certain complexity. Such a model is expected to become the foundation of Artificial General Intelligence (AGI). We have selected several relatively noteworthy points in consciousness research for analysis and discussion, as shown in Appendix D.
@@ 6. Conclusion
In this work, we proposed a generative agent framework based on a computational consciousness structure that uses consciousness channels, protention, memory, and drive to assist in implementing a specific agent. Our agent framework is designed to combine LLMs with mechanisms for synthesizing and retrieving relevant information, thereby providing formatted language model outputs. In our experiment, we deployed an ITCMA in text-based life scenarios and real-world scenarios, both of which effectively handled various tasks and sub-objectives. Our agent performed better than the baseline in the experimental scenarios, demonstrating better utility and the ability to continually evolve to target tasks. In addition, because our framework can be adapted to a specific task without training, it exhibits better flexibility and generalization capabilities. The results of ablation study indicate that each module of ITCM contributes to the utility of the full architecture. Our results indicate that the ITCMA framework has advantages and further potential in handling complex tasks and real-world scenarios.
@@
Acknowledgements.
We thank all all publications support and staff, who wrote and provided helpful comments on previous versions of this document. We gratefully acknowledge the grant from the National Natural Science Fund under Grant 61741206.
@@ Appendix A Phenomenology for the Analysis of the Consciousness Structure
Appendix B Phenomenological Evidence of ITCM
In cognitive science, perception proceeds through three basic forms of development: sensation, perception, and representation. Sensation is the reflection of the individual properties and characteristics of objects in the human brain (with the aid of the body’s nervous system and sensory organs). Perception is the synthesis of various sensations and the reflection of objects as a whole in the human brain. On the basis of perception, a representation is produced. These representations are reproduced through recollection and association and are formed on the basis of multiple perceptions of the same thing or similar things. Representation is the process by which an image of an object that has been perceived is reproduced mentally when the object is not present in front of the individual [34]. In fact, Merleau-Ponty’s [12] study of the perceptual field falls conceptually in the middle of these three forms, and there is a striking similarity between sensation and the elements in the perceptual field, and between representation and the content of the phenomenal field above the perceptual field. This certainly provides some degree of confirmation of Merleau-Ponty’s view of the coherence of consciousness and perception. Does this mean that if we focus only on constructing a model of pseudo-consciousness that can correspond to intelligence, we can turn our attention to first-person phenomenology? We have already introduced this in Appendix A. Here, we primarily present the phenomenological evidence for ITCM.
B.1. Perceptual and Phenomenal Fields
A ”field” is a functional and exclusive activity location of conscious experience, as identified by Husserl [36]. Merleau-Ponty [12] collectively referred to the various experiential ”fields” established by Husserl as ”phenomenal fields.” Merleau-Ponty argues that consciousness works through one part of the phenomenal field, namely, the perceptual field. Only in the perceptual field can silent experience be brought into meaningful expression, completing the actual presentation of consciousness and the conscious activity of the final representation. According to Merleau-Ponty, the perceptual subject is always in the midst of other objects; it is always a part of the field. Each specific experience can only be unified through its own field. Without a mechanism such as the perception field, our perception would not be continuous at all.
What is the content of the perceptual field? According to Husserl’s [36] theory of consciousness, the stream of monads is the most basic unit of meaning in the phenomenal field; a continuous stream of monads forms our perceptions. Libet [37] provides the strongest explanation for this: ”We have shown experimentally that not all nerve cell activities give rise to a conscious experience. For example, a short train of stimulus pulses to sensory cortex elicits responses of many nerve cells without any subjective experience”. A single cell or cell group cannot be the birthplace of conscious experience. Only when a monad becomes part of a monad stream (demonstrating continuous extensibility) can consciousness be constructed, activated, and displayed.
When we obtain these sets of represented monads, the spherical coordinate system that can be transformed by the agent’s actions becomes the agent’s phenomenal field. Visual examples can be used to illustrate some concepts. It is obvious that infants do not reflect on what they see. They take what they see and consider it real. However, an infant’s range of consciousness is not limited to visual stimuli. For objects that are not within their perceptual range, infants with experience will understand the possible changing conditions of these objects [38]. For example, if we often use some ”sleight of hand” tricks to change a piece of candy from the left hand to the right, the infants will eventually go directly to the right hand to obtain the candy.
This implies two situations. The renewal of perception is mandatory in non-reflective consciousness; that is, for the subject, this is ”fact.” In addition, before perception is forced to update, there will still be an impression of the object in the phenomenal field, and it is not static. It will move and transform according to the original experience until perception forces an update.
B.2. Stream of Consciousness and Internal Time-Consciousness
  1. 1.
  2. 2.
  3. 3.
This seems to provide an overly complex structure for every moment of experience in consciousness (i.e., as one extremely long, unbroken event), and retention is complete despite the length of the stream. As time passes, most past objects and experiences quickly lose their distinctiveness and uniqueness; ultimately, they lose any importance that deserves attention.
The components of retention, primal impression, and protention are all field-representations. In the calculation, both retention and primal impression can be obtained, but it seems that we cannot obtain a protention pointing to the future. It should be emphasized that protention is not equivalent to a primal impression at time $t+1$, but through retention and primal impression, consciousness can predict the next moment [39]. Just as we can naturally draw the next segment along a continuous function.
B.3. Memory and Protention
Accidental memory reproduction involves involuntary memory and is built on the similarity between two different sensations at different times. Therefore, the current primal impression is not solely determined by its previous retention. In the current primal impression, a certain memory from the past can be awakened. If it has been awakened at this moment, it is fused into the current consciousness.
B.4. Drive and Emotions
An agent is not just a prediction algorithm. Heidegger [44] regarded plans toward the future as a more fundamental time dimension, thus forming a temporal structure with the future as the priority dimension. This implies that the prediction of protention is not without a trend or is only based on previous trends. There is a driving force in its prediction at time t.
According to Held [45], the source of this drive is “hope.” The hope expressed here is subsumed under the superordinate concept of “expectation.” We believe that the role of expectation is associated with future feelings, and this association is infected by specific emotions. We first expect something and then assign a certain degree of emotion to this expectation.
Of course, hope as a feeling is an indicator of emotion. It is not another short-term impulse but constantly impacts from its own dimension into the action level. Therefore, such a feeling can also be interpreted as a continuous dimensional model.
The consideration of emotions as a driving force here is due to the forced updating of the perceptual field and the prediction of the phenomenal field, which leaves room for the ITCM to introduce emotions. For example, the special feature of the emotion fear is that it is unknown and uncontrollable, which is highly similar to the prediction and perception of objects. The “unknown and uncontrollable” mentioned here corresponds precisely to the dimension of dominance. In addition, for the arousal and pleasure of emotions, we can temporarily assume that the former comes from the amplitude of changes in the monad flow, which is related to passive attention [46], and that the latter comes from satisfying survival needs, such as eating [47]—of course, the actual situation is much more complex.
Appendix C Specific implementation details and examples of ITCMA